OpenStudy (anonymous):

The temperature was 65 degrees at daybreak. Then it dropped two degrees per hour until dusk. If x is the number of hours since daybreak, this equation represents the temperature during the day: y = -2x + 65 OR y = 65 - 2x. Use the equation to predict the temperature six hours later. a)63 b)53 c)61 d)59

OpenStudy (nurali):

y=65-2x put x = 6 y = 65 - 2(6) = 65 -12 =53 degrees

OpenStudy (anonymous):

a

OpenStudy (nurali):

Welcome to Openstudy.

OpenStudy (anonymous):

Thanks! :)

OpenStudy (nurali):

My Pleasure.